OpenStudy (anonymous):

Dimensional Analysis question A very large unit of measurement is the zettameter (zm). 1 zm = 1021 m, but astronomers measures distances in even bigger units called astronomical units (the au). One au is defined to be approximately 107 miles (the average distance between the center of the earth and the center of the sun). An even bigger unit is called the parsec, which is equal to about 2x105 au. When astronomers talk about intergalactic space however it’s even more convenient to use units of kiloparsecs. Given that 1.6 km = 1 mile, about how many kiloparsecs are there in a zettameter?

OpenStudy (anonymous):

Do you understand some of it?

OpenStudy (anonymous):

I got to go in 17min, so the more responsive you are, the faster we finish.

OpenStudy (anonymous):

Yes, the work I've came up with so far is, 10^18 km x (1mi/1.6 km) x (1au/10^7 mi ) x )1 Kiloparsec/2x10^8 au)

OpenStudy (anonymous):

I got 3 x 10^18 Kiloparsecs, but the answer key says the answer is 1/3 x 10^3 Kiloparsecs I don't know what I'm doing wrong

OpenStudy (anonymous):

the conversions should be 1 zm = 10^21 m 1 au = 10^7 mi 1 parsec = 2x10^5 au

OpenStudy (anonymous):

and so 1 kiloparsec= 2x10^8 au

OpenStudy (anonymous):

1000 parsec=1 parsec 1 parsec=2,00,000Au 1Au=107 miles 1zm=1021m

OpenStudy (anonymous):

Yes you right.

OpenStudy (anonymous):

But after calculating, my final answer is different from the answer in the answer key

OpenStudy (anonymous):

one zm has 1021m, so lets start with that.

OpenStudy (anonymous):

is my setup wrong?

OpenStudy (anonymous):

its meant to be 10^21

OpenStudy (anonymous):

I converted 10^21 m to km, which is 10^18 km

OpenStudy (anonymous):

\[\frac{ 1021m }{ 1 } \times \frac{ 0.001km }{ 1m } \times \frac{ 1mile }{ 1.6km } \times \frac{ 1Au }{ 107miles } \times \frac{ 1 parsec }{ 2\times10^5 Au } \times \frac{ 0.001 kiloparsec }{ 1 parsec }\]

OpenStudy (anonymous):

ahh i forgot to convert parsec to kiloparsec thanks so much!!

OpenStudy (anonymous):

Ah, nice job attempting though!

OpenStudy (anonymous):

I have to go now. If you have any questions in 2 hours, I can help.

OpenStudy (anonymous):

alright thanks so much